OpenStudy (anonymous):

Determine the point of inflection, p(t)= 20 / (1+3e^-0.02t)

myininaya (myininaya):

You need to find the 2nd derivative

OpenStudy (anonymous):

yea i need help with that.

myininaya (myininaya):

Ok can you find the first derivative?

OpenStudy (anonymous):

nope.

OpenStudy (amistre64):

just integrate the 2nd dericative to get to the 1st derivative :)

myininaya (myininaya):

Do you know the quotient rule?

OpenStudy (anonymous):

i get confused with the d/dx of the e functions

OpenStudy (anonymous):

yes i do

myininaya (myininaya):

\[\frac{d}{dx}(e^{at})=(at)'e^{at}=ae^{at} \text{ where } a \text{ is a constant }\]

OpenStudy (anonymous):

so d/dx of (1+3e^-0.02t) is: -0.02(3e^-0.02t)?

myininaya (myininaya):

\[p'=\frac{(20)'(1+3e^{-0.02t})-20(1+3e^{-0.02t})'}{(1+3e^{-0.02t})^2}\] Yes!

OpenStudy (anonymous):

so is the first derivative

OpenStudy (anonymous):

0.4(3e^-0.02t)/(1+3e^-0.02t)^2

myininaya (myininaya):

yes

myininaya (myininaya):

Then you can multiply 3 and .4

OpenStudy (anonymous):

okay, now can we go thru the second derivative together?,

OpenStudy (anonymous):

1.2e^-0.02t?

myininaya (myininaya):

Again you will need to use quotient rule

myininaya (myininaya):

You will need the chain rule to for the bottom and for the e thing

OpenStudy (anonymous):

i tried, and its not working

myininaya (myininaya):

\[p''=\frac{(1.2e^{-0.02t})'(1+3e^{-0.02t})^2-(1.2e^{-0.02t})((1+3e^{-0.02t})^2)'}{(1+3e^{-0.02t})^2}\]

OpenStudy (anonymous):

isnt the denominator to the power 4?

myininaya (myininaya):

yes that is a type-o

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

now can we go thru the steps?

myininaya (myininaya):

for the quotient rule?

myininaya (myininaya):

all i did was do [(top)'(bottom)-(top)(bottom)'] /(bottom)^2

OpenStudy (anonymous):

i meant for this question

OpenStudy (anonymous):

so is it? 0.024e^-0.02t(bottom) - top(2)(1+3e^-0.02t)(1.2e^-0.02t)

OpenStudy (anonymous):

oops sorry not 1.2e, i meant 0.06e^...

myininaya (myininaya):

\[(1.2)(-0.02)e^{-0.02t}=(1.2e^{-0.02t})'\]

myininaya (myininaya):

\[2(1+3e^{-0.02t})(0+3(-0.02)e^{-0.02t})=((1+3e^{-0.02t})^2)'\]

myininaya (myininaya):

Just plug in

OpenStudy (anonymous):

this is what i got as my final answer: (0.024e^-0.02t(1-6(1+3e^-0.02t))/ (1+3e^-0.02t)^3